# Staff Scheduling for Demand-Responsive Services

### v1

@@ 1 Introduction
Working time of the employees is a finite and valuable resource in most organizations. Therefore, planning shifts for the staff members in an optimal way is very important in a wide variety of industries, including transportation, hospitality, manufacturing and retail. Especially in settings where the demand for the goods or services provided varies with time, it is very important for the shift plan to ensure that the total deployed workforce at a point of time matches the demand well [thompson1998labor_2]: planning too few employees at a period of high demand leads to lost productivity, planning too many employees during low demand results in an opportunity cost.
The staff scheduling process often needs to take various constraints into account. Some of these constraints may be a consequence of limited availability of a physical resource, such as workspace and equipments [zhang2005equipment, berman1997scheduling]. Some constraints may be legal in nature, e.g. specific amount of breaks between successive shifts, mandatory days off etc. Some constraints may be due to the need to maintain high job satisfaction among the staff [thopsn1999labor_3, gordon2004improving]: Certain staff members may want to work only during the morning or night, for example. Finding any shift plan subject to these constraints me be a difficult task, let alone one maximizing some reward function.
Certain application areas might require specific additional constraints and different objectives. Staff scheduling for transportation services [ernst1998train] often involves spatial constraints on the starting and ending location of the shifts and conformity with a time table of operation. Nurse scheduling in hospitals as well as emergency services need to accommodate different categories of personnel (able to perform different functions), as well as fair distribution of day and night shifts [aickelin2000genetic, mohammadian2019scheduling]. Many industries need to accommodate part time employees [vakharia1992efficient] working only certain days of the week.
Integer programming was introduced as a method of solving staff scheduling problems as early as 1954 [dantzig1954comment], and various integer programming formulations have been proposed since then [keith1979operator, jacobs1996overlapping, bechtold1990implicit, beaumont1997scheduling, brusco2001starting]. The number of decision variables in integer programming models may be very large in realistic settings [glover1986general], leading to very high runtimes. Therefore, various metaheuristics based methods have been studied for solving staff scheduling problems, such as genetic programming [aickelin2000genetic, tanomaru1995staff], tabu search [glover1986general, cavique1999subgraph] and simulated annealing [brusco1993simulated, ernst1998train].
We refer the reader to the review articles [ernst2004survey, alfares2004survey] for comprehensive surveys of the staff scheduling problem and its applications.
@@ 2 Demand Modelling for Staff Scheduling
Staff scheduling in various settings start with a demand modelling step [ernst2004survey, thompson1998labor_2], where for each time step, a required number of active shifts are calculated. We will henceforth refer to this quantity as the desired supply. This is computed by first estimating future demand of the services provided by the organization. Then domain knowledge on how much demand can be served by how much supply is utilized to convert the estimated demand into the desired supply [dantzig1954comment]. In the later stages of the shift scheduling process [ernst2004survey,  p. 6], where shifts are planned while maximizing a certain objective, deviations from the desired supply are penalized. Demand modelling, i.e. converting estimated demand to desired supply can be performed using various approaches [thompson1998labor_2], including the three described below.
@@
We note that both the service standards approach and the economic standards approach assumes that given a supply, i.e. the number of active shifts, and a demand, it is possible to determine the amount of demand that can be served or the amount of reward that can be generated. Crucially, the demand modelling step is usually performed without utilizing any knowledge of the available workforce [thompson1998labor_2].
@@ 4 Problem setting
$x_{i,t}:=\text{ the number of shifts of type }i\text{ {starting} at time }t.$ (1)
@@
Using Equation 2, the resulting mixed-integer convex problem can then be formulated as
@@ 5.3 The reward function
for some $a>0$; see Figure 1 for a visual representation.
$y_{t}$$r_{t}$$d_{t}$$a=4$$a=2$$a=1$
@@ 5.3.1 Demanded rides
see Figure 2 for a visual representation.
@@ 5.4.1 Shift plan must match the number of employed drivers
[…]
Figure 3: Proof of Part 2 of Lemma 1. The dashed rectangle shows a connected component of the bipartite graph $G$. $P_{1}$ can $P_{2}$ can be swapped, keeping the graph bipartite, because there is no edge between the connected component and the rest of the graph.
We note that the proof of Lemma 1 is constructive, and therefore, we can use it to assign the shifts to the drivers in a feasible way.
@@ 5.5 The Full Program
s.t. $y_{t}=\sum_{\tau=t-\delta+1}^{t}x_{\tau},$ for all $\text{for all }t\in[T],$ (16b)
$y_{t}\leq c,$ for all $\text{for all }t\in[T],$ (16c)
@@
$z_{t}=\sum_{\tau=t-\delta-\beta+1}^{t}x_{t},$ for all $\text{for all }t\in[T],$ (16e)
$z_{t}\leq N,$ for all $\text{for all }t\in[T],$ (16f)
$x_{t}\geq 0$ for all $\text{for all }t\in[T],$ (16g)
$x_{t}\in\mathbb{Z}$ for all $\text{for all }t\in[T],$ (16h)
@@ 5.6 Benchmarking against shift agnostic optimum
$\frac{\partial f_{t}(y^{*}_{t})}{\partial y_{t}}+\mu_{t}$ $=\lambda,$ for all $\text{for all }t\in[T]$ (20)
@@
The KKT conditions [boyd2004convex] guarantee that any solution to the maximization problem (19) must satisfy (20).
@@ 5.7 Benchmarking against approaches using a separate demand modelling step
Here we outline a method to compare the quality of the shift plans generated by the process described above with shift plans generated using a traditional demand modelling step [thompson1998labor_2]. As described in Section 2, these approaches first convert the demand to required working staff at each point of time. At the second step, the shift plans are generated by solving a mixed-integer convex problem that minimizes the sum of quadratic deviations from the required working staff at each point of time.
@@
s.t. $y_{t}=\sum_{\tau=t-\delta+1}^{t}x_{\tau},$ for all $\text{for all }t\in[T],$ (27b)
$y_{t}\leq c,$ for all $\text{for all }t\in[T],$ (27c)
@@
$z_{t}=\sum_{\tau=t-\delta-\beta+1}^{t}x_{t},$ for all $\text{for all }t\in[T],$ (27e)
$z_{t}\leq N,$ for all $\text{for all }t\in[T],$ (27f)
$x_{t}\geq 0$ for all $\text{for all }t\in[T],$ (27g)
$x_{t}\in\mathbb{Z}$ for all $\text{for all }t\in[T].$ (27h)
@@ 6 Results
Now we present the results of solving the mixed-integer convex program described in Section 5.5, using the open source SCIP optimization suite [BestuzhevaEtal2021OO]. To this end, we reduce the problem to a mixed-integer linear program (MIP) by first replacing the reward function $f_{t}$ defined in (17) with a piecewise linear concave approximation
@@ 6.1 Solving the MIP to produce a shift plan
In the following, we will use the term shift agnostic optimum for both the optimal value as well as the optimal solution of the optimization problem (19), whenever it is clear from the context which one is meant. Figure 4 shows a solution to the MIP formulation (29) for $N=10$, $s=5$, $\delta=8$, $\beta=8$, $d_{\max}=10$, and $a=2$. The solution is reasonably close to the shift agnostic optimum, following the periodic demand pattern. However, the various shift constraints prevent the solution from perfectly matching the shift agnostic optimum. In particular, the solution does not have the smooth shape of the shift agnostic optimum, but is piecewise constant. Moreover, the solution does not reach the peak of the shift agnostic optimum in the middle of the week. In the following sections, we will investigate the impact of the various shift constraints on the quality of the solution and how relaxing them helps to overcome these limitations.
@@ 6.2.1 Impact of the total number of drivers
Since every started shift is active for $\delta$ time steps, the active shifts curve $\textbf{y}=(y_{t})_{t\in[T]}$ is the sum of $sN$ step functions. For a small number of drivers $N$, it is therefore difficult to approximate the smooth shift agnostic optimum $\textbf{y}^{*}=(y^{*}_{t})_{t\in[T]}$. As the number of drivers increases, the approximation becomes more accurate, and the gap between the solution y and the shift agnostic optimum $\textbf{y}^{*}$ decreases. Figure 5 illustrates this effect. For an increasing number of drivers $N$, we computed the corresponding optimal shift plan x and its supply curve y as well as the shift agnostic optimum $\textbf{y}^{*}$. For comparability, we normalized each supply curve by dividing it by the total working hours $sN\delta$.
For each number of drivers, we have computed the served trips using (8), and also the served trips for the shift agnostic optimum. The gap between the two is then the lost trips due to shift restrictions.These lost trips, divided by the total served trips due to the shift agnostic optimum, i.e. the relative gap from optimum supply as defined in (22), are plotted in panel (b) of Figure 5. We observe that the higher the number of drivers, the smoother the supply curve of the optimal shift plan and the smaller the relative gap from optimum supply. Notice, however, that even for as many as 200 drivers, the shift agnostic optimum is not reached. The reason for this is that, since every driver needs to work $s$ shifts per week, only $1/s$ of the total number of shifts can be active at the same time. Thus, the high peak in the middle of the week cannot be completely fulfilled. This limitation can be overcome by reducing the number of shifts per driver, as we will see in the next section.
@@ 6.2.2 Impact of weekly shifts per driver
If the total working time $sN\delta$ and the shift length $\delta$ is fixed but both the number of drivers $N$ and the number of shifts per driver $s$ are flexible, then it is beneficial to choose more drivers with fewer shifts; see Figure 6. As an intuitive example, consider two scenarios, one with $N$ drivers with $s$ shifts each, and another with $sN$ drivers with only one shift each. Then any shift plan for the first scenario can also be realized in the second scenario by assigning each of the $s$ shifts of a driver in the first scenario to $s$ different drivers in the second scenario. Therefore, the optimal shift plan in the second scenario is at least as good as the optimal shift plan in the first scenario. Moreover, while in the first scenario only $N$ shifts can be active at the same time, in the second scenario $sN$ shifts can be active at the same time, allowing to reach higher peaks.
@@ 6.2.3 Impact of shift lengths
Finally, if the total working time $sN\delta$ and the number of shifts per driver $s$ is fixed but both the shift length $\delta$ and the number of drivers $N$ are flexible, then it is beneficial to choose more drivers with shorter shifts. In this way, we have more step functions with smaller support that can approximate the smooth shift agnostic optimum better; see Figure 7. Consider, for example, five different choices of shift length, $\delta\in\left\{1,2,4,8,16\right\}$. Then any solution with a shift length of $\delta=2^{i}$, $i>0$, and $N$ drivers can also be realized by $2N$ drivers with shift length $\delta^{\prime}=\delta/2$ by substituting each shift of length $\delta$ with two shifts of length $\delta^{\prime}$. Therefore, the smaller $i$ is, the better the resulting solution.
@@ 6.3 Comparison to approaches with a separate demand modelling step
Figure 8: Our shift planning method achieves higher total reward than approaches with a separate demand modelling step. We have plotted the relative gap from optimum supply (21) for different values of the number of drivers $N$. All three approaches achieve smaller relative gap from optimum supply as the number of drivers increases, but our approach achieves the least relative gap from optimum supply for all values of $N$. For service standards, the parameter $c$ was set to $0.8$, and for economic standards, the parameter $c$ was set to $1$. Other parameters were set to $s=5$, $\delta=8$, $\beta=9$, $d_{\max}=0.75N$, $a=2$.
@@
We also see in Figure 9 that the quality of the shift plans is highly dependent on the choice of service parameter in (24) for the service standards approach, and on the choice of the cost parameter in (25) for the economic standards approach. This demonstrates another advantage of our approach of directly maximizing the total reward, namely that it is not dependent on any such internal parameters.
@@ 7 Outlook
First of all, it would be interesting to apply our approach to scenarios with more than one shift type. This would require generalizing the constraints (12) and Lemma 1 to ensure that the optimized shift plan is rosterable. Also, we have so far not considered breaks within a shift, but this can be easily accommodated by redefining active shifts in (2).
@@
In this article, we have limited ourselves to producing a shift plan without assigning the shifts to individual employees. For only one shift type, as we have considered in Section 5, the proof of Lemma 1 can readily be turned into an algorithm for shift assignment. A generalization of Lemma 1 and an algorithm for shift assignment to multiple shift types is left for future research. We note that a different approach for assigning shifts to employees is to redefine shift types as described in Section 4 so that each shift type corresponds to shifts from a single employee. However, this comes at the cost of a larger optimization problem.
@@ 9 Declaration of interests
[…]

### v2

@@ 1 Introduction
Working time of the employees is a finite and valuable resource in most organizations. Therefore, planning shifts for the staff members in an optimal way is very important in a wide variety of industries, including transportation, hospitality, manufacturing and retail. Especially in settings where the demand for the goods or services provided varies with time, it is very important for the shift plan to ensure that the total deployed workforce at a point of time matches the demand well [undefs]: planning too few employees at a period of high demand leads to lost productivity, planning too many employees during low demand results in an opportunity cost.
The staff scheduling process often needs to take various constraints into account. Some of these constraints may be a consequence of limited availability of a physical resource, such as workspace and equipments [undefv, undefd]. Some constraints may be legal in nature, e.g. specific amount of breaks between successive shifts, mandatory days off etc. Some constraints may be due to the need to maintain high job satisfaction among the staff [undeft, undefn]: Certain staff members may want to work only during the morning or night, for example. Finding any shift plan subject to these constraints me be a difficult task, let alone one maximizing some reward function.
Certain application areas might require specific additional constraints and different objectives. Staff scheduling for transportation services [undefk] often involves spatial constraints on the starting and ending location of the shifts and conformity with a time table of operation. Nurse scheduling in hospitals as well as emergency services need to accommodate different categories of personnel (able to perform different functions), as well as fair distribution of day and night shifts [undef, undefq]. Many industries need to accommodate part time employees [undefu] working only certain days of the week.
Integer programming was introduced as a method of solving staff scheduling problems as early as 1954 [undefj], and various integer programming formulations have been proposed since then [undefp, undefo, undefc, undefb, undefh]. The number of decision variables in integer programming models may be very large in realistic settings [undefm], leading to very high runtimes. Therefore, various metaheuristics based methods have been studied for solving staff scheduling problems, such as genetic programming [undef, undefr], tabu search [undefm, undefi] and simulated annealing [undefg, undefk].
We refer the reader to the review articles [undefl, undefa] for comprehensive surveys of the staff scheduling problem and its applications.
@@ 2 Demand Modelling for Staff Scheduling
Staff scheduling in various settings start with a demand modelling step [undefl, undefs], where for each time step, a required number of active shifts are calculated. We will henceforth refer to this quantity as the desired supply. This is computed by first estimating future demand of the services provided by the organization. Then domain knowledge on how much demand can be served by how much supply is utilized to convert the estimated demand into the desired supply [undefj]. In the later stages of the shift scheduling process [undefl,  p. 6], where shifts are planned while maximizing a certain objective, deviations from the desired supply are penalized. Demand modelling, i.e. converting estimated demand to desired supply can be performed using various approaches [undefs], including the three described below.
@@
We note that both the service standards approach and the economic standards approach assumes that given a supply, i.e. the number of active shifts, and a demand, it is possible to determine the amount of demand that can be served or the amount of reward that can be generated. Crucially, the demand modelling step is usually performed without utilizing any knowledge of the available workforce [undefs].
@@ 4 Problem setting
[…]
@@
Using Equation˜2, the resulting mixed-integer convex problem can then be formulated as
@@ 5.3 The reward function
for some $a>0$; see Figure˜1 for a visual representation.
$y_{t}$$r_{t}$$d_{t}$a=4a=4a=2a=2a=1a=1
@@ 5.3.1 Demanded rides
see Figure˜2 for a visual representation.
@@ 5.4.1 Shift plan must match the number of employed drivers
$S_{1}$$S_{2}$$P_{1}$$P_{2}$
Figure 3: Proof of Part 2 of Lemma˜1. The dashed rectangle shows a connected component of the bipartite graph $G$. $P_{1}$ can $P_{2}$ can be swapped, keeping the graph bipartite, because there is no edge between the connected component and the rest of the graph.
We note that the proof of Lemma˜1 is constructive, and therefore, we can use it to assign the shifts to the drivers in a feasible way.
@@ 5.5 The Full Program
s.t. $y_{t}=\sum_{\tau=t-\delta+1}^{t}x_{\tau},$ $\text{for all }t\in[T],$ (16b)
$y_{t}\leq c,$ $\text{for all }t\in[T],$ (16c)
@@
$z_{t}=\sum_{\tau=t-\delta-\beta+1}^{t}x_{t},$ $\text{for all }t\in[T],$ (16e)
$z_{t}\leq N,$ $\text{for all }t\in[T],$ (16f)
$x_{t}\geq 0$ $\text{for all }t\in[T],$ (16g)
$x_{t}\in\mathbb{Z}$ $\text{for all }t\in[T],$ (16h)
@@ 5.6 Benchmarking against shift agnostic optimum
$\frac{\partial f_{t}(y^{*}_{t})}{\partial y_{t}}+\mu_{t}$ $=\lambda,$ $\text{for all }t\in[T]$ (20)
@@
The KKT conditions [undeff] guarantee that any solution to the maximization problem (19) must satisfy (20).
@@ 5.7 Benchmarking against approaches using a separate demand modelling step
Here we outline a method to compare the quality of the shift plans generated by the process described above with shift plans generated using a traditional demand modelling step [undefs]. As described in Section˜2, these approaches first convert the demand to required working staff at each point of time. At the second step, the shift plans are generated by solving a mixed-integer convex problem that minimizes the sum of quadratic deviations from the required working staff at each point of time.
@@
s.t. $y_{t}=\sum_{\tau=t-\delta+1}^{t}x_{\tau},$ $\text{for all }t\in[T],$ (27b)
$y_{t}\leq c,$ $\text{for all }t\in[T],$ (27c)
@@
$z_{t}=\sum_{\tau=t-\delta-\beta+1}^{t}x_{t},$ $\text{for all }t\in[T],$ (27e)
$z_{t}\leq N,$ $\text{for all }t\in[T],$ (27f)
$x_{t}\geq 0$ $\text{for all }t\in[T],$ (27g)
$x_{t}\in\mathbb{Z}$ $\text{for all }t\in[T].$ (27h)
@@ 6 Results
Now we present the results of solving the mixed-integer convex program described in Section 5.5, using the open source SCIP optimization suite [undefe]. To this end, we reduce the problem to a mixed-integer linear program (MIP) by first replacing the reward function $f_{t}$ defined in (17) with a piecewise linear concave approximation
@@ 6.1 Solving the MIP to produce a shift plan
In the following, we will use the term shift agnostic optimum for both the optimal value as well as the optimal solution of the optimization problem (19), whenever it is clear from the context which one is meant. Figure˜4 shows a solution to the MIP formulation (29) for $N=10$, $s=5$, $\delta=8$, $\beta=8$, $d_{\max}=10$, and $a=2$. The solution is reasonably close to the shift agnostic optimum, following the periodic demand pattern. However, the various shift constraints prevent the solution from perfectly matching the shift agnostic optimum. In particular, the solution does not have the smooth shape of the shift agnostic optimum, but is piecewise constant. Moreover, the solution does not reach the peak of the shift agnostic optimum in the middle of the week. In the following sections, we will investigate the impact of the various shift constraints on the quality of the solution and how relaxing them helps to overcome these limitations.
@@ 6.2.1 Impact of the total number of drivers
Since every started shift is active for $\delta$ time steps, the active shifts curve $\textbf{y}=(y_{t})_{t\in[T]}$ is the sum of $sN$ step functions. For a small number of drivers $N$, it is therefore difficult to approximate the smooth shift agnostic optimum $\textbf{y}^{*}=(y^{*}_{t})_{t\in[T]}$. As the number of drivers increases, the approximation becomes more accurate, and the gap between the solution y and the shift agnostic optimum $\textbf{y}^{*}$ decreases. Figure˜5 illustrates this effect. For an increasing number of drivers $N$, we computed the corresponding optimal shift plan x and its supply curve y as well as the shift agnostic optimum $\textbf{y}^{*}$. For comparability, we normalized each supply curve by dividing it by the total working hours $sN\delta$.
For each number of drivers, we have computed the served trips using (8), and also the served trips for the shift agnostic optimum. The gap between the two is then the lost trips due to shift restrictions.These lost trips, divided by the total served trips due to the shift agnostic optimum, i.e. the relative gap from optimum supply as defined in (22), are plotted in panel (b) of Figure˜5. We observe that the higher the number of drivers, the smoother the supply curve of the optimal shift plan and the smaller the relative gap from optimum supply. Notice, however, that even for as many as 200 drivers, the shift agnostic optimum is not reached. The reason for this is that, since every driver needs to work $s$ shifts per week, only $1/s$ of the total number of shifts can be active at the same time. Thus, the high peak in the middle of the week cannot be completely fulfilled. This limitation can be overcome by reducing the number of shifts per driver, as we will see in the next section.
@@ 6.2.2 Impact of weekly shifts per driver
If the total working time $sN\delta$ and the shift length $\delta$ is fixed but both the number of drivers $N$ and the number of shifts per driver $s$ are flexible, then it is beneficial to choose more drivers with fewer shifts; see Figure˜6. As an intuitive example, consider two scenarios, one with $N$ drivers with $s$ shifts each, and another with $sN$ drivers with only one shift each. Then any shift plan for the first scenario can also be realized in the second scenario by assigning each of the $s$ shifts of a driver in the first scenario to $s$ different drivers in the second scenario. Therefore, the optimal shift plan in the second scenario is at least as good as the optimal shift plan in the first scenario. Moreover, while in the first scenario only $N$ shifts can be active at the same time, in the second scenario $sN$ shifts can be active at the same time, allowing to reach higher peaks.
@@ 6.2.3 Impact of shift lengths
Finally, if the total working time $sN\delta$ and the number of shifts per driver $s$ is fixed but both the shift length $\delta$ and the number of drivers $N$ are flexible, then it is beneficial to choose more drivers with shorter shifts. In this way, we have more step functions with smaller support that can approximate the smooth shift agnostic optimum better; see Figure˜7. Consider, for example, five different choices of shift length, $\delta\in\left\{1,2,4,8,16\right\}$. Then any solution with a shift length of $\delta=2^{i}$, $i>0$, and $N$ drivers can also be realized by $2N$ drivers with shift length $\delta^{\prime}=\delta/2$ by substituting each shift of length $\delta$ with two shifts of length $\delta^{\prime}$. Therefore, the smaller $i$ is, the better the resulting solution.
@@ 6.3 Comparison to approaches with a separate demand modelling step
Figure 8: Our shift planning method achieves higher total reward than approaches with a separate demand modelling step. We have plotted the relative gap from optimum supply (21) for different values of the number of drivers NN. All three approaches achieve smaller relative gap from optimum supply as the number of drivers increases, but our approach achieves the least relative gap from optimum supply for all values of NN. For service standards, the parameter $c$ was set to $0.8$, and for economic standards, the parameter $c$ was set to 11. Other parameters were set to $s=5$, $\delta=8$, $\beta=9$, $d_{\max}=0.75N$, $a=2$.
@@
We also see in Figure˜9 that the quality of the shift plans is highly dependent on the choice of service parameter in (24) for the service standards approach, and on the choice of the cost parameter in (25) for the economic standards approach. This demonstrates another advantage of our approach of directly maximizing the total reward, namely that it is not dependent on any such internal parameters.
@@ 7 Outlook
First of all, it would be interesting to apply our approach to scenarios with more than one shift type. This would require generalizing the constraints (12) and Lemma˜1 to ensure that the optimized shift plan is rosterable. Also, we have so far not considered breaks within a shift, but this can be easily accommodated by redefining active shifts in (2).
@@
In this article, we have limited ourselves to producing a shift plan without assigning the shifts to individual employees. For only one shift type, as we have considered in Section˜5, the proof of Lemma˜1 can readily be turned into an algorithm for shift assignment. A generalization of Lemma˜1 and an algorithm for shift assignment to multiple shift types is left for future research. We note that a different approach for assigning shifts to employees is to redefine shift types as described in Section˜4 so that each shift type corresponds to shifts from a single employee. However, this comes at the cost of a larger optimization problem.
@@ 9 Declaration of interests
Acknowledgements
We thank our colleagues at MOIA GmbH and MOIA Operations Germany GmbH for their valuable input, and especially Yılmaz Arslanoğlu, who co‑developed the initial prototype of the employee-scheduling solution presented in this paper.